# Capturing System Drift with Time Series Calibration for Global 21-cm Cosmology Experiments

### v1

@@ Abstract
To achieve the sensitivity required to detect signals from neutral hydrogen from the Cosmic Dawn and Epoch of Reionisation it is critical to have a well-calibrated instrument which has a stable calibration over the course of the observation. Previous calibration methods do not explicitly use the time information available and make assumptions on the impedance matching of the reference sources. Here we present a new calibration method based on noise wave parameters which fits a calibration solution over time and frequency to the data, interpolating the solutions to the times at which the antenna is being measured. To test this method we simulate a dataset using measurements of the REACH receiver, modelling a low noise amplifier which is drifting over time. Fitting a polynomial surface in frequency and time to the simulated data demonstrates that we can remove the drift in the calibrated solution over time but leaves a chromatic residual. We further show that we can remove assumptions on the reflection coefficients of the reference noise source and the cold load, reducing degeneracies in the parameter fits. Applying this new calibration equation and surface fitting method to the simulated data removes the chromatic residual in the calibrated spectrum and recovers the parameters to within 0.06% of the truth and a 97% reduction in the RMSE of the spectrum of the validation source compared with previous calibration methods. For two parameters we report up to six times smaller fit error after the degeneracies are removed from the time-based calibration.
@@ 1 Introduction
As the global 21-cm signal absorption depth is predicted to be small – between 6868\,mK and 201201\,mK (Dhandha et al., 2025) – while the free-free and synchrotron emission from the galactic foregrounds will be of order $10^{3}-10^{4}\,$K (Shaver et al., 1999), we require highly sensitive experiments to separate the signal from the foregrounds. To do so requires long integration times, with a REACH-like experiment requiring approximately 6.5 hours of integration at minimum to reduce the radiometer noise down to the level at which it can detect the EDGES signal (de Lera Acedo et al., 2022). As a result of this long integration time, we need to maintain a stable calibration over the observation to achieve the sensitivity necessary to detect the 21-cm signal. Previous calibration methods (Rogers & Bowman, 2012; Roque et al., 2021) have not considered calibration over time and so require the instrument to be kept stable over the course of the observation.
@@ 2.2 Removing Impedance Matching Assumptions and Degeneracies
Equation 1 makes the assumption that the noise source and cold load are impedance matched, i.e. $\Gamma_{\mathrm{NS}}=\Gamma_{\mathrm{L}}$ [^1] Note that equation 1 does not make the assumption that the noise source and cold load are impedance matched to the LNA.. In reality, where the impendance matching is imperfect, Roque et al. (2021) accounts for the mismatch by fitting frequency-dependent functions for $T_{\mathrm{NS}}^{\prime}$ and $T_{\mathrm{L}}^{\prime}$ which capture the realistic system [^2] Similarly, in the EDGES calibration this is accounted for by introducing two parameters $C_{1}$ and $C_{2}$ (Monsalve et al., 2017). These two parameters also fit for the path difference between their noise source and the antenna.. However, this introduces degeneracies into your noise wave parameter fit as it can be shown that
@@ 2.4 Surface Fitting
Previous REACH noise wave calibration methods (Roque et al., 2021, 2025; Kirkham et al., 2025) do not explicitly consider the time information when calibrating the receiver. This can be problematic since the antenna is measured in between measurements of the calibrators and so, should the state of the LNA change significantly during an observation, we would require an interpolation of the noise wave parameters to accurately calibrate the antenna. The times at which sources are measured in a example REACH observation is plotted in figure 1 where the twelve calibrators are measured then the antenna spectra is measured five times.

### v2

@@ Abstract
To achieve the sensitivity required to detect signals from neutral hydrogen from the Cosmic Dawn and Epoch of Reionisation it is critical to have a well-calibrated instrument which has a stable calibration over the course of the observation. Previous Bayesian calibration methods do not explicitly use the time information available and make assumptions on the impedance matching of the reference sources. Here we present a new calibration method based on noise wave parameters which fits a calibration solution over time and frequency to the data, interpolating the solutions to the times at which the antenna is being measured. To test this method we simulate a dataset using measurements of the REACH receiver, modelling a low noise amplifier which is drifting over time. Fitting a polynomial surface in frequency and time to the simulated data demonstrates that we can remove the drift in the calibrated solution over time but leaves a chromatic residual. We further show that we can remove assumptions on the reflection coefficients of the reference noise source and the cold load, reducing degeneracies in the parameter fits. Applying this new calibration equation and surface fitting method to the simulated data removes the chromatic residual in the calibrated spectrum and recovers the parameters to within 0.06% of the truth and a 97% reduction in the RMSE of the spectrum of the validation source compared with previous calibration methods. For two parameters we report up to six times smaller fit error after the degeneracies are removed from the time-based calibration.
@@ 1 Introduction
As the global 21-cm signal absorption depth is predicted to be small – with an amplitude $\lesssim 200\,$mK (Dhandha et al., 2025) – while the free-free and synchrotron emission from the galactic foregrounds will be of order $10^{3}-10^{4}\,$K (Shaver et al., 1999), we require highly sensitive experiments to separate the signal from the foregrounds. To do so requires long integration times, with a REACH-like experiment requiring approximately 6.5 hours of integration at minimum to reduce the radiometer noise down to the level at which it can detect the EDGES signal (de Lera Acedo et al., 2022). As a result of this long integration time, we need to maintain a stable calibration over the observation to achieve the sensitivity necessary to detect the 21-cm signal. Previous calibration methods (Rogers & Bowman, 2012; Roque et al., 2021) have not considered calibration over time and so require the instrument to be kept stable over the course of the observation.
@@ 2.2 Removing Impedance Matching Assumptions and Degeneracies
[…]
@@ 2.4 Surface Fitting
Previous REACH noise wave calibration methods (Roque et al., 2021, 2025; Kirkham et al., 2025) do not explicitly consider the time information when calibrating the receiver [^2] Another time-based calibration method can be seen in Leeney et al. (2025) which adds time as an input to the calibration neural network.. This can be problematic since the antenna is measured in between measurements of the calibrators and so, should the state of the LNA change significantly during an observation, we would require an interpolation of the noise wave parameters to accurately calibrate the antenna. The times at which sources are measured in a example REACH observation is plotted in figure 1 where the twelve calibrators are measured then the antenna spectra is measured five times.